OpenStudy (anonymous):

Find the value of: 2sqrt(3/25)

OpenStudy (anonymous):

0.346...

OpenStudy (anonymous):

\[\frac{2}{5}\sqrt3\]

OpenStudy (mertsj):

\[2\sqrt{\frac{3}{25}}=2\frac{\sqrt{3}}{\sqrt{25}}=2\frac{\sqrt{3}}{5}\]

OpenStudy (anonymous):

Thank you ver much!!! :)

OpenStudy (mertsj):

yw